Consider the system of linear equations below.
3x+y=-4
-5x+y=8
A. Graph the system of equations on the coordinate plane. Label each line on your graph and show any work you do.
B. Use the graph form Part A to estimate the solution for the system of equations. Explain how you determined your estimate.

Answer by mananth(16946)

3 x + 1 y = -4 .............1
-5 x + 1 y = 8 .............2

Find the y & x intercepts of both lines and draw the lines.
The point of intersection of the lines is the solution of the system
equation1
when x= 0 y= -4 ( 0 , -4 )
when y = 0 x= -1.33 ( -1.33 , 0 )
Equation 2
when x= 0 y= 8 ( 0 , 8 )
when y = 0 x= -1.6 ( -1.6 , 0 )

The solution is (-1.5,0.5)